function g(x) = -1/x. Determine the average rate of change of the function between x = 3 and x = 3 + h.
I really am struggling with this and another problem. But the rest of my problems I flew through easily. I feel I am having problems with the (3+h) part of this problem, because when x= "3" or just "h" I get it.

Answer by stanbon(75887)

function g(x) = -1/x. Determine the average rate of change of the function between x = 3 and x = 3 + h.
------
Average = [g(3+h)-g(3)]/[3+h-3]
-----
= [-1/(3+h)-(-1/3)]/[h]
----
= [-1/(3+h)+(1/3)]/h
----
= [(-3+3+h)/(3(3+h))]/h
----------
= [h/(9+3h)]/h
--
= 1/(9+3h)
=================
Cheers,
Stan H.
=================